Assume you started with 0.252 g of unknown iron compound, and got a yield of 0.191 g Fe2O3. What's the percent of iron in the unknown?
Answer #1

Molar mass of Fe2O3 is 159.7 g/mol.
0.191 g of Fe2O3 corresponds to \dfrac {0.191}{159.7} = 0.001196 moles. They will be obtained from 2 \times 0.001196 = 0.002392 moles of Fe.
The atomic mass of Fe is 55.8 g/mol.
0.002392 moles of Fe corresponds to 55.8 \times 0.002392 = 0.1335 g
The percent of iron in the unknown = \dfrac {0.1335 \times 100}{0.191} = 69.9 %